(A-Grade) If 8% of all people in a certain region are unemployed, find the probability that in a sample of 200 people, there are fewer than 10 people who are unemployed. (approximate Binomial by Normal Distribution)
Answer #1

Solution:

Given that,

p = 8% = 0.08

1 - p = 1 - 0.08 = 0.92

n = 200

Here, X\sim BIN (200 , 0.08)

then,

n*p = 200 * 0.08 = 16

n(1- p) = 200 * 0.92 = 184

Since both are greater than 5 , we can use normal approximation to binomial.

According to normal approximation binomial,

X \rightarrow Normal

Mean = \mu = n*p = 16

SD = \sigma =\sqrt{}n*p*(1-p) = \sqrt{} [200 * 0.08 * 0.92]  = 3.83666521865

P(X < 10)

= P((x - \mu ) / \sigma < (10 - 16)/3.83666521865)

= P(z < -1.56)

Probability = 0.0594